# Learning Low-Dimensional Embeddings for Black-Box Optimization

### v1

@@ 4 Numerical examples
We evaluate the performance of Meta-GLIS on two numerical examples. The first involves the Rosenbrock function, a well-established benchmark in optimization. The second focuses on automated calibration of MPC hyperparameters. Computations are carried out on a server equipped with an NVIDIA RTX 3090 GPU. To guarantee reproducibility of the results, all Python scripts are made available in the GitHub repository https://github.com/buswayne/meta-glis.
4.1 Rosenbrock benchmark
In the multidimensional case, the Rosenbrock function to be minimized is defined as:
@@ 4.1 Rosenbrock benchmark
where $\theta_{1}\in\mathbb{R},\theta_{2}\in\mathbb{R},\theta_{3}\in\mathbb{R}^{n_{x}-1}$ are random realization sampled, respectively, from the following uniform distributions $\mathcal{U}_{[10,1000]},\mathcal{U}_{[0.1,10]},(\mathcal{U}_{[0.1,10]},\ldots,\mathcal{U}_{[0.1,10]})$. For the nominal case, $\theta=\left[100,1,\boldsymbol{1}\right]$, the function is non-convex and has a narrow, curved valley containing the global minimum at $(x_{1}^{\star},\ldots,x_{n_{x}}^{\star})=(1,\ldots,1)$.
The meta-dataset is obtained from $N=500$ instances of the parameter $\theta$, sampled from the uniform distributions described above, with $n_{x}=20$. We solved each problem with differential evolution [13] with 1000 generations, and setting bounds $\mathcal{X}=[-2.5,2.5]^{20}$ on the input domain. We then created a meta-dataset $\mathcal{D}=\left\{\theta^{(i)},\{x^{(i)}_{k},f^{(i)}_{k}\}_{k=1}^{1000}\right\}_{i=1}^{500}$. Similarly, we generated a separate test dataset $\mathcal{D}^{\mathrm{test}}=\left\{\theta^{(i)},\{x^{(i)}_{k},f^{(i)}_{k}\}_{k=1}^{1000}\right\}_{i=1}^{100}$ to evaluate the reconstruction capabilities of an autoencoder with latent dimension $n_{z}=3$, when trained as discussed in Section 3.2. The encoder and decoder are fully connected neural networks, each with two hidden layers. The encoder has $n_{x}$ inputs, a first hidden layer with 128 units, a second hidden layer with 64 units, and $n_{z}$ outputs. The decoder has a symmetric structure. The time required to train the autoencoder is $1229$ s (approximately 20.4 minutes).
@@
[…]
Results are reported in Fig. 2, which shows the best value of the Rosenbrock objective function achieved at each iteration of the GLIS and Meta-GLIS algorithm. The figure presents both the average and standard deviation over the 100 Monte Carlo runs. The results highlight the advantages of Meta-GLIS, which achieves near-optimal solutions after approximately 20 iterations, whereas GLIS remains far from the global optimum even after 100 iterations. This underscores the well-known advantages of running GLIS, and more in general BBO, in low-dimensional spaces.
A further analysis is provided in Fig. 3, which illustrates a single instance of the Monte Carlo runs. The figure shows the points sampled by GLIS and Meta-GLIS over the algorithm’s iterations, projected onto the same planes as in Fig. 1. It is evident that in GLIS, which directly optimizes over $x$, the sampled points (blue circles) after 100 iterations almost uniformly cover the search space $\mathcal{X}$. In contrast, in Meta-GLIS, where optimization is performed over a lower-dimensional space, the sampled points (orange circles) remain within the learned manifold by construction, and the solution (orange stars) converges much closer to the optimal value.
Another advantage of Meta-GLIS over GLIS is the time required to optimize the acquisition function: on average, 0.027 s for Meta-GLIS vs. 0.149 s for GLIS. This is because the acquisition function in Meta-GLIS is defined over a lower-dimensional space ($n_{z}=3$) compared to GLIS ($n_{x}=20$). As for the time required to fit the surrogate model, it is similar for both Meta-GLIS and GLIS (approximately 0.002 s), as it primarily depends on the number of sampled points rather than the input dimensionality.
Figure 2: Comparison of the average performance obtained with GLIS (blue) and Meta-GLIS (red) over $N^{\mathrm{test}}=100$ runs. Global optimum (green). Shaded areas correspond to standard deviation over $100$ Monte Carlo runs.
@@ 4.2 Automated calibration of MPC hyperparameters
We consider the same problem in [3], which uses GLIS for automatically tuning of the hyperparameters of a model predictive controller. The system to be controlled is a cart-pole described by the differential equations:
$(M+m)\ddot{p}+mL\ddot{\psi}-mL\dot{\psi}^{2}\sin{\psi}+b\dot{p}$ $=F$ (9)
@@
with $p$, $\psi$ and $F$ being, respectively, the cart position (m), the pendulum angle (rad), and the input force on the cart (N). The force $F$ is regulated by the MPC to track a position reference $p^{\mathrm{ref}}$ (dashed red line in Fig. 4), while maintaining the pole in a vertical position (i.e., $\psi\approx 0$) for a pre-defined duration of the experiment $T^{\mathrm{exp}}$. In addition, whenever: $(i)$ the pendulum falls ($|\psi|>\frac{\pi}{6}$), $(ii)$ the cart position exceeds the track length ($|p|>1.1$ m) or $(iii)$ numerical errors occur in solving the MPC problem, the controller is stopped at a time $T^{\mathrm{stop}}<T^{\mathrm{exp}}$.
@@
$J_{\rm MPC}=$ $\sum\nolimits_{k=0}^{N_{p}-1}(y_{t+k|t}-y^{\text{ref}}_{t+k})^{T}Q_{y}(y_{t+k|t}-y^{\text{ref}}_{t+k})+$
$+$ $\sum\nolimits_{k=0}^{N_{p}-1}(u_{t+k|t}-u^{\text{ref}}_{t+k})^{T}Q_{u}(u_{t+k|t}-u^{\text{ref}}_{t+k})+$
$+$ $\sum\nolimits_{k=0}^{N_{p}-1}\Delta u_{t+k|t}^{T}Q_{\Delta u}\Delta u_{t+k|t}+Q_{\epsilon}\eta^{2},$ (11)
subject to input and output constraints (see [3] for details on this case study). In (4.2), $y=[p,\ \psi]$; $u=F$; $N_{p}$ is the prediction horizon; $u^{\text{ref}}_{t}$ and $y^{\text{ref}}_{t}$ are the input and output reference; respectively, at a generic time step $t$; $\eta$ is a slack variable used in [3] to soften the input and output constraints; and $\Delta u_{t}$ is the input variation.
@@
Figure 5: Empirical distributions of the optimal solutions over $100$ test system realizations (solid lines) and corresponding $3000$ queries sampled during the optimization (shaded areas) for each MPC hyperparameter. GLIS (blue); Meta-GLIS with nonlinear (orange) and linear decoder (purple).
The hyper-parameters calibration problem is formalized as the minimization of the following closed-loop performance index over the entire duration $T_{\mathrm{exp}}$ of the experiment:
$\tilde{J}^{\mathrm{cl}}=\ln\left(\int_{t=0}^{T_{\mathrm{exp}}}10|p^{\mathrm{ref}}(t)-p(t)|+30|\phi(t)|\;dt\right)+$
@@
To generate the meta-dataset $\mathcal{D}$, we sample the parameters $\theta=[M,m,b,f_{\varphi},L]^{\top}$ of the cart-pole dynamics equations from distributions $[\mathcal{U}_{[0.5,2.5]},\mathcal{U}_{[0.2,1.0]},\mathcal{U}_{[0.1,0.5]},\allowbreak\mathcal{U}_{[0.1,0.5]},\mathcal{U}_{[0.3,1.5]}]$, generating a total of $N=200$ system configurations. In the generation of the meta-dataset, the cost function (4.2) is minimized with GLIS over $500$ iterations, leading to a meta-dataset $\mathcal{D}=\left\{\theta^{(i)},\{x^{(i)}_{k},f^{(i)}_{k}\}_{k=1}^{500}\right\}_{i=1}^{200}$ with $000200\times 500=100^{\prime}$ 000 points. These explored points are used for training an autoencoder with the same architecture as the previous example ($n_{z}=3$), with $\alpha=0.5$.
@@
For a fair comparison, Meta-GLIS is evaluated against standard GLIS that also leverages information from the meta-dataset $\mathcal{D}$. In particular, we initialize GLIS with the optimal hyperparameter vector associated to the central parameter $\theta_{\mathrm{o}}=[1.5,0.6,0.3,0.3,0.9]$.

### v2

@@ 3.3 Meta-GLIS: GLIS over a low-dimensional space
4 Sub-optimality quantification
The goal of this section is to establish probabilistic theoretical bounds that quantify the gap between solving the original problem (1) and the reduced-dimensional problem (2) over the latent set $\mathcal{Z}$.
4.1 Definitions
For a given parameter $\theta$ taking values in a measurable set $\Theta\subseteq\mathbb{R}^{n_{\theta}}$ and a random variable $\xi\in\Xi\subseteq\mathbb{R}^{n_{\xi}}$, let $\mathcal{A}_{x}:\Theta\times\Xi\to\mathcal{X}$ be a deterministic map from $(\theta,\xi)$ to a value $x=\mathcal{A}_{x}(\theta,\xi)$, $x\in\mathcal{X}$. In our work, $\mathcal{A}_{x}$ represents a numerical optimization algorithm (e.g., gradient descent, particle swarm optimization, Bayesian optimization), $\xi$ the random realizations characterizing the algorithm (e.g., initial conditions, mini-batches in stochastic gradient descent, etc.), and $\mathcal{A}_{x}(\theta,\xi)$ the solution returned by $\mathcal{A}_{x}$ when applied to problem (1) with parameters $\theta$ and for random algorithm’s parameters $\xi$. Alternatively, $\mathcal{A}_{x}$ can be the mapping that returns a global minimizer of problem (1). In this case, $x^{\star}(\theta)=\mathcal{A}_{x}(\theta,\xi)$ depends only on the parameters $\theta$ characterizing the objective function ff [^1] In case of multiple global optima, $x^{\star}(\theta)$ may also depend on some random variables $\xi$ if $x^{\star}(\theta)$ is selected at random from the set of global optimizers..
Analogously, let $\mathcal{A}_{z}:\Theta\times\Omega\to\mathcal{Z}$ be a deterministic mapping such that $\mathcal{A}_{z}(\theta,\omega)$ represents the solution of the reduced-dimensional problem (2) returned by algorithm $\mathcal{A}_{z}$, with $\omega\in\Omega$ capturing the random settings involved in the reduced-space optimization procedure. Note that $\mathcal{A}_{x}$ and $\mathcal{A}_{z}$ do not need to be the same numerical algorithms.
We measure the relative gap in performance when optimizing in the reduced space via the relative performance gap defined as:
$\Psi(\theta,\xi,\omega):=\frac{f(\mathrm{D}(\mathcal{A}_{z}(\theta,\omega));\theta)-f(\mathcal{A}_{x}(\theta,\xi);\theta)}{|f(\mathcal{A}_{x}(\theta,\xi);\theta)|+\epsilon},$ (9)
where $\epsilon>0$ is a small constant introduced to avoid possible divisions by zero. In the special case
$\mathcal{A}_{x}(\theta,\xi)=x^{\star}(\theta),\qquad\mathcal{A}_{z}(\theta,\omega)=z^{\star}(\theta),$ (10)
the performance gap $\Psi(\theta,\xi,\omega)$ measures the theoretical loss between the optimal solution of problem (1) and that of the reduced-dimensional problem (2). However, this value can be computed only if problems (1) and (2) can be solved exactly at the global optimum. More generally, when $\mathcal{A}_{x}$ and $\mathcal{A}_{z}$ denote numerical optimization algorithms, the performance gap may also take negative values as the global optimum of problem (1) might not be reached by the algorithm $\mathcal{A}_{x}$.
Let $p(\xi)$ and $p(\omega)$ be distributions over the random variables $\xi$ and $\omega$, and $\{(\theta^{(i)},\allowbreak\xi^{(i)},\allowbreak\omega^{(i)})\}_{i=1}^{m}$ be a validation set of i.i.d. draws from product measure $p(\theta)\otimes p(\xi)\otimes p(\omega)$. The validation set thus corresponds to $m$ realizations of the performance gap $\Psi$, namely, with $i=1,\dots,m$,
$f^{(i)}:=f(x(\theta^{(i)},\xi^{(i)});\theta^{(i)}),\quad\Psi^{(i)}:=\Psi(\theta^{(i)},\xi^{(i)},\omega^{(i)}).$ (11)
Finally, let $(\theta^{\rm new},\xi^{\rm new},\omega^{\rm new})$ be an independent draw from $p(\theta)\otimes p(\xi)\otimes p(\omega)$, not used during training or validation, with corresponding objective value $f^{\rm new}$ and performance gap $\Psi^{\rm new}$.
We denote by $\mathbb{P}$ the underlying probability measure induced by the product distribution $p(\theta)\otimes p(\xi)\otimes p(\omega)$. Accordingly, for any measurable set $A\subseteq\mathbb{R}$, we write
$\mathbb{P}_{\Psi}(A):=\mathbb{P}(\Psi\in A)=\mathbb{P}\left(\Psi^{-1}(A)\right)$ (12)
to denote the probability distribution of the random variable $\Psi$. Similarly, let $S=\{\Psi^{(i)}\}_{i=1}^{m}$ denote the validation set. We write $\mathbb{P}_{S}$ for the probability distribution of $S$, induced by the i.i.d. draws $\{(\theta^{(i)},\xi^{(i)},\omega^{(i)})\}_{i=1}^{m}\sim p(\theta)\otimes p(\xi)\otimes p(\omega)$.
4.2 Generalization bounds on the performance gap
Theorem 1 (Generalization bounds)
For any $\alpha,\delta\in(0,1)$, let $k^{\star}:=\Big\lceil m\,(1-\alpha+\varepsilon_{m})\Big\rceil$ and $\varepsilon_{m}:=\sqrt{\tfrac{\log(2/\delta)}{2m}}$. Assume that the number of validation samples $m$ is large enough so that $\alpha\geq\varepsilon_{m}$ (or equivalently, $k^{\star}\leq m$). Then, with probability at least $1-\delta$ over the validation set $S=\left\{\Psi^{(i)}\right\}_{i=1}^{m}$, we have
$\mathbb{P}_{\Psi}\!\left(\Psi\leq\Psi_{(k^{\star})}\right)\;\geq\;1-\alpha,$ (13)
where $\Psi_{(k)}$ denotes the $k$-th order statistic of the validation values, i.e., if
$\Psi_{(1)}\leq\Psi_{(2)}\leq\dots\leq\Psi_{(m)},$ (14)
then $\Psi_{(k)}$ is the $k$-th smallest value in the ordered sequence. Equivalently,
$\mathbb{P}_{S}\Big(\,\mathbb{P}_{\Psi}\big(\Psi\leq\Psi_{(k^{\star})}\,\big|\,S\big)\;\geq\;1-\alpha\,\Big)\;\geq\;1-\delta.$ (15)
$\blacksquare$
The proof of the theorem is reported in the Appendix.
Theorem 1 has the following interpretation. To construct the validation set, we generate $m$ i.i.d. realizations $\{(\theta^{(i)},\xi^{(i)},\omega^{(i)})\}_{i=1}^{m}$ from the distribution product $p(\theta)\otimes p(\xi)\otimes p(\omega)$. For each triple $(\theta^{(i)},\xi^{(i)},\omega^{(i)})$, we run the two optimization algorithms $\mathcal{A}_{x}$ in the original space (e.g., PSO) and $\mathcal{A}_{z}$ in the reduced latent space (e.g., GLIS), and we compute the corresponding performance gap
$\Psi^{(i)}:=\Psi(\theta^{(i)},\xi^{(i)},\omega^{(i)}),\qquad i=1,\dots,m.$ (16)
The theorem guarantees that, with probability at least $1-\delta$ over the draw of the validation set $S$, the order statistic $\Psi_{(k^{\star})}$ provides an upper bound on the $(1-\alpha)$-quantile of the true distribution of $\Psi$. Equivalently, for a new independent draw $(\theta^{\rm new},\xi^{\rm new},\omega^{\rm new})\sim p(\theta)\otimes p(\xi)\otimes p(\omega)$ and corresponding performance gap $\Psi^{\rm new}=\Psi(\theta^{\rm new},\xi^{\rm new},\omega^{\rm new})$, we have
$\mathbb{P}_{\Psi}\!\left(\Psi^{\rm new}\;\leq\;\Psi_{(k^{\star})}\,\Big|\,S\right)\;\geq\;1-\alpha,$ (17)
and this condition holds with probability at least $1-\delta$ with respect to the randomness of the validation set $S$.
@@ 5 Numerical examples
We evaluate the performance of Meta-GLIS on two numerical examples. The first involves the Rosenbrock function [13], a well-established benchmark problem in nonlinear optimization. The second focuses on the automated calibration of a set of MPC tuning knobs. Computations are carried out on a server equipped with an NVIDIA RTX 3090 GPU. To guarantee reproducibility of the results, all Python scripts are made available in the GitHub repository https://github.com/buswayne/meta-glis.
5.1 Rosenbrock function benchmark
The Rosenbrock function to be minimized is defined as:
@@ 5.1 Rosenbrock function benchmark
where $\theta_{1}\in\mathbb{R},\theta_{2}\in\mathbb{R},\theta_{3}\in\mathbb{R}^{n_{x}-1}$ are random realization sampled, respectively, from the following uniform distributions $\mathcal{U}_{[10,1000]},\mathcal{U}_{[0.1,10]},(\mathcal{U}_{[0.1,10]},\ldots,\mathcal{U}_{[0.1,10]})$. For the nominal case, $\theta=\left[100,1,\boldsymbol{1}\right]$, where $\boldsymbol{1}\in\mathbb{R}^{n_{x}}$ is a vector with all entries equal to 1, the function is non-convex and has a narrow, curved valley containing the global minimum at $(x_{1}^{\star},\ldots,x_{n_{x}}^{\star})=(1,\ldots,1)$.
The meta-dataset is obtained from $N=500$ instances of the parameter $\theta$, sampled from the uniform distributions described above, with $n_{x}=20$. We solved each problem by employing a differential evolution algorithm [18] with 1000 generations over the feasible input domain $\mathcal{X}=[-2.5,2.5]^{20}$. We then created a meta-dataset $\mathcal{D}=\left\{\theta^{(i)},\{x^{(i)}_{k},f^{(i)}_{k}\}_{k=1}^{1000}\right\}_{i=1}^{500}$. Similarly, we generated a separate test dataset $\mathcal{D}^{\mathrm{test}}=\left\{\theta^{(i)},\{x^{(i)}_{k},f^{(i)}_{k}\}_{k=1}^{1000}\right\}_{i=1}^{100}$ to evaluate the reconstruction capabilities of an autoencoder with latent dimension $n_{z}=3$, when trained as discussed in Section 3.2. The encoder and decoder are fully connected neural networks, each with two hidden layers. The encoder has $n_{x}$ inputs, a first hidden layer with 128 units, a second hidden layer with 64 units, and $n_{z}$ outputs. The decoder has a symmetric structure. The time required to train the autoencoder is approximately 20 minutes.
@@
The Meta-GLIS algorithm is applied in the learned $n_{z}=3$-dimensional latent space, with $M_{\rm init}=2n_{z}=6$ initial random points generated using LHS within the box $\mathcal{Z}=[0,\ 1]^{3}$. For comparison, optimization is also performed using GLIS over the original $20$-dimensional input domain. In this case, $2n_{x}=40$ initial points are randomly generated within the box $2.5,\ 2.5]^{20}$ using LHS. A Monte Carlo analysis is conducted on 100 newly generated problem instances that were not used for training the autoencoder.
The obtained numerical results are reported in Fig. 2, which shows the best value of the Rosenbrock objective function achieved at each iteration of the GLIS and Meta-GLIS algorithm. The figure presents both the average and standard deviation over 100 Monte Carlo runs. The results highlight the advantages of Meta-GLIS, which achieves near-optimal solutions after approximately 20 iterations, whereas GLIS remains far from the global optimum even after 100 iterations. This underscores the well-known advantages of running GLIS, and more in general BBO, in low-dimensional spaces, due to the reduced curse of dimensionality caused by exploration.
Figure 2: Comparison of the average performance obtained with GLIS (blue) and Meta-GLIS (red) over $N^{\mathrm{test}}=100$ runs. Global optimum (green). Shaded areas correspond $\pm$ one standard deviation over $100$ Monte Carlo runs.
@@
This advantage can be formally quantified by applying the theoretical guarantees of Theorem 1. We generated $M=1000$ validation instances of the optimization problem and solved them using PSO, GLIS, and Meta-GLIS, which allowed us to compute the corresponding probabilistic bounds on the performance gap $\Psi$.
For $\delta=0.05$ and $\alpha=0.1$, the computed bound on the performance gap between Meta-GLIS and PSO is $1.14\%$ (blue dashed line in Fig. 4(a)). Thus, with probability at least $95\%$ (i.e., $1-\delta$), the likelihood that a new instance optimized with Meta-GLIS leads to a performance gap w.r.t. PSO lower than $1.14\%$ is at least $90\%$ (i.e., $1-\alpha$). This result is consistent with the empirical $90$-th percentile on the test set (orange dashed line), which gives a smaller gap of $0.68\%$, thereby validating the bound.
By applying the same reasoning to compare standard GLIS (operating on the full space $x$) against Meta-GLIS, we obtain the result shown in Fig. 4(b). In this case, Meta-GLIS demonstrates a clear advantage: $90\%$ of the test instances achieve at least a 68.87% relative improvement when optimization is performed in the latent space. This result can qualitatively be appreciated also in Fig. 3, which illustrates a single instance of the Monte Carlo runs. The figure shows the points sampled by GLIS and Meta-GLIS over the algorithm’s iterations, projected onto three planes as in Fig. 1. It is evident that in GLIS, which directly optimizes over $x$, the sampled points (blue circles) after 100 iterations almost uniformly cover the search space $\mathcal{X}$. In contrast, in Meta-GLIS, where optimization is performed over a lower-dimensional space, the sampled points (orange circles) remain within the learned manifold by construction, and the solution (orange stars) converges much closer to the optimal value.
Another advantage of Meta-GLIS over GLIS is the time required to optimize the acquisition function: on average, 0.027 s for Meta-GLIS vs. 0.149 s for GLIS. This is because the acquisition function in Meta-GLIS is defined over a lower-dimensional space ($n_{z}=3$) compared to GLIS ($n_{x}=20$). As for the time required to fit the surrogate model, it is similar for both Meta-GLIS and GLIS (approximately 0.002 s), as it primarily depends on the number of sampled points rather than on the input dimensionality.
(a) Meta-GLIS vs. PSO
(b) Meta-GLIS vs. GLIS
Figure 4: Empirical CDFs of the performance gap using Meta-GLIS compared with PSO (panel (a)) and GLIS (panel (b)) over $100$ test instances. Theoretical bound from Theorem 1 (blue line); empirical 90%-quantile on the test set (orange line).
@@ 5.2 Automated calibration of MPC hyperparameters
As another case study, we consider the same problem in [3], which uses GLIS for automatically tuning certain hyperparameters of a model predictive controller. The system to be controlled is a cart-pole described by the differential equations:
$(M+m)\ddot{p}+mL\ddot{\psi}-mL\dot{\psi}^{2}\sin{\psi}+b\dot{p}$ $=F,$ (19)
@@
with $p$, $\psi$ and $F$ being, respectively, the cart position (m), the pendulum angle (rad), and the input force on the cart (N). The force $F$ is regulated by the MPC to track a position reference $p^{\mathrm{ref}}$ (dashed red line in Fig. 7), while maintaining the pole in a vertical position (i.e., $\psi\approx 0$) for a pre-defined duration of the experiment $T^{\mathrm{exp}}$. In addition, whenever: $(i)$ the pendulum falls ($|\psi|>\frac{\pi}{6}$), $(ii)$ the cart position exceeds the track length ($|p|>1.1$ m), or $(iii)$ numerical errors occur in solving the MPC problem, the controller is stopped at a time $T^{\mathrm{stop}}<T^{\mathrm{exp}}$.
@@
$J_{\rm MPC}=$ $\displaystyle{\sum_{k=0}^{N_{p}-1}}\Big((y_{t+k|t}-y^{\text{ref}}_{t+k})^{T}Q_{y}(y_{t+k|t}-y^{\text{ref}}_{t+k})$
$+(u_{t+k|t}-u^{\text{ref}}_{t+k})^{T}Q_{u}(u_{t+k|t}-u^{\text{ref}}_{t+k})$
$+\Delta u_{t+k|t}^{T}Q_{\Delta u}\Delta u_{t+k|t}\Big)+Q_{\epsilon}\eta^{2},$ (21)
subject to input and output constraints (see [3] for details). In (5.2), $y=[p,\ \psi]$; $u=F$; $N_{p}$ is the prediction horizon; $u^{\text{ref}}_{t}$ and $y^{\text{ref}}_{t}$ are the input and output reference; respectively, at a generic time step $t$; $\eta$ is a slack variable used in [3] to soften the input and output constraints; and $\Delta u_{t}$ is the input variation.
@@
Figure 5: Empirical distributions of the optimal solutions over $100$ test system realizations (solid lines) and corresponding $3000$ queries sampled during the optimization (shaded areas) for each MPC hyperparameter: GLIS (blue); Meta-GLIS with nonlinear (orange) and linear decoder (purple).
The hyperparameter-calibration problem is formalized as the minimization of the following closed-loop performance index over the entire duration $T_{\mathrm{exp}}$ of the experiment:
$\tilde{J}^{\mathrm{cl}}=\ln\left(\int_{t=0}^{T_{\mathrm{exp}}}10|p^{\mathrm{ref}}(t)-p(t)|+30|\phi(t)|\;dt\right)$
@@
To generate the meta-dataset $\mathcal{D}$, we sample the parameters $\theta=[M,m,b,f_{\varphi},L]^{\top}$ of the cart-pole dynamics equations from the uniform distribution over $\mathcal{U}_{[0.5,2.5]}\times\mathcal{U}_{[0.2,1.0]}\times\mathcal{U}_{[0.1,0.5]}\times\mathcal{U}_{[0.1,0.5]}\times\mathcal{U}_{[0.3,1.5]}$, generating a total of $N=200$ system configurations. In the generation of $\mathcal{D}$, the cost function (5.2) is minimized with GLIS over $500$ iterations, leading to $\mathcal{D}=\left\{\theta^{(i)},\{x^{(i)}_{k},f^{(i)}_{k}\}_{k=1}^{500}\right\}_{i=1}^{200}$ with $200\times 500=100\,000$ points. These explored points are used for training an autoencoder with the same architecture as the previous example ($n_{z}=3$), with $\lambda=0.5$.
@@
For a fair comparison, Meta-GLIS is evaluated against standard GLIS that also leverages information from the meta-dataset $\mathcal{D}$. In particular, we initialize GLIS with the optimal hyperparameter vector associated to the central parameter $\theta_{\mathrm{o}}=[1.5,\allowbreak 0.6,\allowbreak 0.3,\allowbreak 0.3,\allowbreak 0.9]$.